# Upper bound on the multiplicity of rational and Du Bois singularities

### v1

@@ 1. Introduction
Based on Theorem 1.1, we propose the following multiplicity bound in terms of minimal log discrepancies.
@@
Here, $\mathrm{mld}_{x}(X)$ denotes the minimal log discrepancy of $X$ at $x$, i.e. the infimum of the log discrepancies $a(E;X)$ over all divisors $E$ with center $x$. Theorem 1.1 confirms the case $k=0$, and a minor refinement of its proof yields the case $k=1$. The equality statement is known in low dimensions: for surfaces with klt singularities by Artin [Artin66]*Corollary 6, and for threefolds with terminal singularities by Kakimi [Kakimi]*Theorem 2.1 and by Takagi–Watanabe [TW04]*Proposition 3.10.

### v2

@@ 1. Introduction
Motivated by Theorem 1.1 and Shibata [Shibata17a]*Proposition 3.9, we propose the following multiplicity bound in terms of minimal log discrepancies.
@@
Here, $\mathrm{mld}_{x}(X)$ denotes the minimal log discrepancy of $X$ at $x$, i.e. the infimum of the log discrepancies $a(E;X)$ over all divisors $E$ with center $x$. Note that [Shibata17a] proves the statement when $X$ is klt and $\mathrm{mld}_{x}(X)$ is replaced by the log canonical threshold of the maximal ideal at $x\in X$. Theorem 1.1 confirms the case $k=0$, and a minor refinement of its proof yields the case $k=1$. The equality statement is known in low dimensions: for surfaces with klt singularities by Artin [Artin66]*Corollary 6, and for threefolds with terminal singularities by Kakimi [Kakimi]*Theorem 2.1 and by Takagi-Watanabe [TW04]*Proposition 3.10.